OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

terenzreignz (terenzreignz):

First step in this sort of word problem is to represent: Assign a values where needed.

OpenStudy (anonymous):

Thats the thing idk where they go

terenzreignz (terenzreignz):

Ok, let's assign values to their present ages, then. If we let Tim's age be x, what would be Coach's age?

OpenStudy (anonymous):

if the kid is ten and the coach is 20 the kid would be 13

OpenStudy (anonymous):

Let's make Coach Johnson J and Tim T T + 10 = J T - 3 = 50% or something like that haha @allietreetz don't just give the answer geez

OpenStudy (anonymous):

wooow no need to dis people i was only try n 2 help

OpenStudy (anonymous):

why did she use 50% ?

terenzreignz (terenzreignz):

I honestly don't know :)

OpenStudy (anonymous):

lol oops I should've done x2 for twice his age...meh bad

terenzreignz (terenzreignz):

But uhmm... If Tim's age is x, then Coach's age would be x + 10, right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

no wiat...

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Oh god lol

OpenStudy (anonymous):

xD I'm just gonna like...leave now xD

terenzreignz (terenzreignz):

So, if Tim's present age is x, and Coach's present age is x + 10, what would be their ages three years ago? :)

OpenStudy (anonymous):

Let age of Johnson is x years and that of Tim is y years.. So, according to first sentence of problem : x = y + 10 -------------1 Now, three years ago, Coach was (x - 3) and Tim was (y - 3) (x - 3) = 2(y - 3) x - 3 = 2y - 6 x = 2y - 3 -----------------2 Equate 1 and 2 equations, y + 10 = 2y - 3 y = 10 + 3 = 13.. So, x = y + 10 = 13 + 10 = 23 So the age of Tim (y) is 13 years...

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

^ The Winner!